OpenStudy (anonymous):

what is the slope between the following points (4,-1) and (-1,4)

OpenStudy (anonymous):

slope=y2-y1/x2-x1 =(4-(-1))/(-1-4) =(4+1)/(-5) = 5/(-5) =-1